THE PRODUCT OF 8/11 AND A NUMBER IS 11/3; WHAT IS THE OTHER NUMBER?
Answer by checkley75(3666)

8/11*X=11/3 CROSS MULTIPLY THUS:
24X=121
X=121/24
X=5 1/24
PROOF
8/11*121/24=11/3
8*121/11*24=11/3
968/264=11/3 CROSS MULTIPLY
2904=2904